### v1

@@ 1 Introduction
Evaluating the text generated by large language models (LLMs) has long been a challenging task, Traditional manual evaluation methods are not only time-consuming and laborious but also expensive[2]. Although methods like BLEU[14], Rouge[10], and METEOR[1] have achieved success in scenarios such as machine translation, these automated evaluation methods are limited in the context of open-ended text generation[11].
@@
Recently, LLMs have been used as evaluators such as G-Eval[4], but these methods exhibit unstable and uncertain evaluation effects [15][16]. Even certain collaboration frameworks could alleviate this problem by employing multi-agent discussion, e.g. ChatEval [3], however, the current methods of multi-agent collaboration remains limited to simple interactions, without fully harnessing the potential for agents’ thinking and planning. Additionally, reaching a consensus within multi-agent discussion frameworks continues to be a challenging issue. Furthermore, traditional text evaluation models typically provide only a score without explaining it, making it difficult for reviewers to trust the reliability of these scores. They still need to manually identify errors, obviously slowing down the collection of bad cases and, consequently, affecting the pace of model iteration in industrial scenarios.
@@
To address the above challenges, this paper introduces a Multi-Agent Text Evalution Framework (MATEval). In this framework, we simulate the human collaborative process in evaluating texts generated by LLMs and propose a novel multi-agent discussion strategy. This strategy integrates self-reflection[13] and Chain-of-Thought (CoT)[17] concepts, as self-reflection focuses on understanding the depth of issues but may lead to rigid thinking. Meanwhile, strategies based on the CoT emphasize the refinement of problems but may lack in-depth analysis of specific issues. So we combine the two approaches by guiding agents through prompts to decompose evaluation questions and focus on only one sub-question in each discussion round. During each round of the discussion, agents engage in self-reflection, considering peer inputs to enrich issue comprehension. This approach strengthens agents’ self-assessment and critical thinking, broadening their evaluation scope for open-ended text and aligning results more closely with human evaluations.
@@
Furthermore, our framework introduces a feedback mechanism at the end of each discussion round to evaluate the quality and efficiency of the discussions, encouraging agents to reach consensus. The comprehensive evaluation report generated by our framework details error types, specific locations, in-depth explanations, and corresponding scores. For practical applications in industry, we provide two report formats: a question-and-answer format for strategy analysis and a text report format for business personnel for aiding error identification quickly and LLMs’ iterative improvement. Our framework has achieved significant results in the story text evaluation task at Alipay, markedly enhancing the efficiency of model iteration.
@@
    We propose a novel method to integrate self-reflection and CoT in multi-agent framework. Additionally, we creatively introduce a feedback mechanism at the end of each discussion round to resolve disagreements and facilitate the achievement of consensus.
@@
    We conduct comprehensive experiments on two English and two Chinese story text datasets, including one constructed based on Alipay’s business story text dataset. And our experimental results showcase the effectiveness of our framework and its high correlation with human evaluations.  [^1] We have made the datasets and results used in our experiments publicly available at https://anonymous.4open.science/r/MATEval-EA92. Due to user privacy of Alipay, we cannot make the “Ant” dataset public.
@@ 2 Related Work
LLM-based Evaluators: GPTScore[4] utilizes models such as GPT-3 to evaluate text quality, predicated on the assumption that generative pre-trained models assign higher probabilities to high-quality generated texts in accordance with given instructions and context. Recent studies also explore the potential of using ChatGPT as an NLG evaluator, where ChatGPT directly scores texts based on prompts[15]. G-Eval[12] demonstrates the evaluation of NLG outputs using prompts in LLMs like ChatGPT through Chain-of-Thought (CoT) methods.
@@ 3.1 Design of the Framework
The framework accepts text as input, which may contain various quality issues. The output is a detailed evaluation report outlining error type, location, explanation, and score. We present the results in two formats: one is a Q&A format that is conducive to evaluating correlation, so that we can easily extract correlation scores to calculate similarity. The other is a report format that is conducive to iteration by relevant business personnel. They can directly understand text issues and iterate models by reading analysis reports, thereby improving efficiency. This is shown in the right part of Figure 1.
@@ 3.2 Application of Different Roles
Figure 2: The diagram including prompts and dialogue that incorporates a process of discussion with self-reflection, CoT, feedback mechanisms and final summary.
Evaluator Agent: The core element in the framework is the evaluator, for which we use GPT-4 to conduct multi-round evaluations and responses which guided through carefully designed prompts. The evaluator stores and processes statements from other agents, using this as a reference for dialogue history. Each agent not only receives responses from others but also provides their own statements, with the entire process requiring minimal human intervention.
@@
Summarizer Agent: After all discussions are concluded, the summarizer compiles the entire process and outcomes. It provides a Q&A format evaluation report, detailing the identification, analysis, and scoring of various issues. Additionally, we provide a comprehensive text-based format evaluation report that includes detailed problem descriptions and are easy to read to help improve model performance in industrial production.
@@ 3.4 Combined Self-reflection and CoT
Combined Self-reflection and CoT: As shown in Algorithm 1. Combining self-reflection and the CoT is an important strategy employed in our framework. Agents autonomously decompose the question according to the prompt, focusing on one sub-question in each round of multiple discussions: $\mathcal{D}(\mathcal{Q})=\{\mathcal{Q}_{\text{1}},\mathcal{Q}_{\text{2}},\ldots,\mathcal{Q}_{\text{n}}\}$, $\mathcal{P}_{\text{i}}=\mathcal{I}(\mathcal{Q}_{\text{i}},\mathcal{H})$ , where $\mathcal{I}$ represents the formation of preliminary ideas based on the prompt, $\mathcal{Q}$ is the evaluation task, $\mathcal{Q}_{\text{n}}$ is the sub-question and $\mathcal{H}$ is the history information. Then optimize their statements through self-reflection: $\mathcal{R}_{\text{i}}=\mathcal{S}(\mathcal{P}_{\text{i}},\mathcal{H})$. Next, update the history: =$\mathcal{H}=\mathcal{H}\cup\{\mathcal{R}_{i}\}$. After each round of discussion, a feedback provider evaluates the discussion to reduce repetition and disagreement: $\mathcal{E}=\text{E}(\mathcal{H})$ Finally, a summarizer compiles all statements to produce the evaluation report: $\mathcal{R}_{\text{final}}=\text{Summary}(\mathcal{H})$. The overall prompt and the flow of the discussion are illustrated in Figure 2.
@@ 4.2 Dataset
We mainly apply our framework to the evaluation of story texts generated by LLMs in Alipay business scenarios. So in the experiment, we mainly focus on two open-ended story datasets: ROCStories (ROC)[6] and WritingPrompts (WP)[6]. Considering GPT-4’s context length limitations and the need for storing multi-round discussion contexts, we truncate WP stories to the first 200 words, ensuring textual integrity at the truncation point. To test model generalizability, we conduct similar experiments on two Chinese datasets. These include the Chinese LOng Text understanding and generation (LOT)[7] dataset, based on hundreds of words from human-written Chinese stories, and a dataset of Chinese fairy tales constructed using prompts from Alipay’s business data with GPT-3.5 named Ant, mainly involving fables and fairy tales. Considering GPT-4’s request rate limits and high usage costs, we select the first 200 stories from each dataset for multi-agent discussion experiments.
@@ 4.3 Compared Methods
Referenced Metrics: The BLEU[14] score is used to evaluate lexical similarity between candidate and reference texts. ROUGE-L[10] focuses on the longest common subsequence to assess the fluency and coherence of texts. And RUBER-BERT[5], an enhancement of the original RUBER model with BERT’s contextual embeddings, includes both referenced and unreferenced versions. The referenced version, RUBER-BERTr, measures similarity between candidate responses and reference texts using BERT word embeddings.
@@
  • Chain-of-Thought (CoT): Agent breaks down the assessment problem through prompts, solving one sub-problem in each discussion round.
@@ 4.4 Experimental Results
Our experiments on the ROC and WP datasets are presented in Tables 1, using MATEval framework strategies to evaluate narrative texts. We calculate Spearman ($\rho$) and Kendall ($\tau$) correlation coefficients to compare model’s evaluations with human judgments.
Table 1: Correlation of evaluation results with human judgment using different models and different strategies of MATEval on the ROC/WP dataset, where SA stands for Single-Agent, SR denotes Self-Reflection, and CoT represents Chain-of-Thought. The symbols $\rho$/$\tau$ respectively indicate Spearman/Kendall correlation. The highest correlation values are highlighted in bold.
@@
In analyzing the strategies used within the MATEval framework, we found that employing self-reflection or CoT independently produce unstable results across different error types, In some cases, these methods even underperformed single-agent evaluations. This might be due to inherent flaws when applying these strategies separately. For example, self-reflection can lead to rigid thinking in multiple discussion rounds, where agents often repeat earlier content without new insights. On the other hand, using CoT alone often results in superficial and divergent perspectives, leading to inefficient analyses. And CoT method typically offers only a basic analysis of each error type without delving deeper.
The combination of self-reflection and CoT achieved the best overall correlation, particularly excelling in evaluating Logical Inconsistency, Discontinuity, and Inappropriate Lexical Choice. It significantly improves the evaluation of Discontinuity compared to single-agent method, demonstrating the framework’s high sensitivity to textual coherence. However, its effectiveness was lower for Repetition and Factual Error. Agents often misidentified emotionally similar sentences as repetitive, despite clear definitions. Furthermore, the framework’s evaluation of Factual Errors was limited by LLMs constraints, specifically the absence of external knowledge affecting common sense error detection, highlighting the need to integrate external knowledge for future multi-agent framework enhancements.
@@ 4.5 Ablation Study
To assess the effectiveness of different modules in MATEval, we conducted ablation experiments on the ROC dataset, involving the removal of the feedback mechanism, omitting Q&A format explanations, and not using a multi-agent approach. Table 3 demonstrate that the complete MATEval framework surpasses its ablated versions, confirming the significance of both feedback mechanisms, explanations and multi-agent methods. Specifically, this establishes the importance of our feedback mechanisms in promoting discussion consensus and enhancing relevance. It also proves that providing scoring explanations by LLMs significantly enhances evaluation effectiveness.
@@ 4.6 Generalization Experiments
To verify the generalizability of our framework across different languages and in the industrial field, we conducted experiments on two Chinese datasets: LOT and a story text dataset Ant, constructed based on business data from Alipay. The findings in Table 2 were similar to those from English datasets. Interestingly, in Chinese, single-agent evaluations sometimes show better results. This might be due to the unique aspects of Chinese language and sentence structure. It indicates that for different languages, the setup and optimization of agents might need custom adjustments to achieve the best performance.
Figure 3: The schematic diagram illustrating the comparison of results generated by the MATEval framework and other methods when applied to a story text example in the industrial field.
@@ 4.7 Case Study
To demonstrate the effectiveness of the MATEval framework in practical industrial applications, we selected an example of a story text for illustration, which presented in Table 3. In this example, we compared our framework with manual evaluation, single-agent evaluation, and the One_by_One evaluation method without any strategy. The experimental table show that our method maintains basic consistency with human evaluation, while the other methods exhibit some omissions, thus illustrating the high correlation between our method and human evaluation.
@@ 5 Conclusion
In this paper, we proposed the MATEval framework, which enhances the evaluation performance of open-ended story text generated by LLMs in the industrial field. Extensive experiments show that MATEval’s evaluation results on two classic story datasets that are more aligned with human preferences compared to existing methods. In the Alipay industrial scenario, our framework significantly improves review efficiency and evaluation accuracy, serving as an effective aid. [^2] This work was supported by Ant Group Research Fund

### v2

@@ 1 Introduction
Evaluating the text generated by large language models (LLMs) has long been a challenging task, Traditional manual evaluation methods are not only time-consuming and laborious but also expensive[2]. Although methods like BLEU[14], Rouge[10], and METEOR[1] have achieved success in scenarios such as machine translation, these automated evaluation methods are limited in the context of open-ended text generation[11]. Recently, LLMs have been used as evaluators such as G-Eval[4], but these methods exhibit unstable and uncertain evaluation effects [15][16]. Even certain collaboration frameworks could alleviate this problem by employing multi-agent discussion, e.g. ChatEval [3], however, the current methods of multi-agent collaboration remain limited to simple interactions, without fully harnessing the potential for agents’ thinking and planning. Additionally, reaching a consensus within multi-agent discussion frameworks continues to be a challenging issue. Furthermore, traditional text evaluation models typically provide only a score without explaining it, making it difficult for reviewers to trust the reliability of these scores. They still need to manually identify errors, obviously slowing down the collection of bad cases and, consequently, affecting the pace of model iteration in industrial scenarios.
@@
To address the above challenges, this paper introduces a Multi-Agent Text Evaluation Framework (MATEval). In this framework, we simulate the human collaborative process in evaluating texts generated by LLMs and propose a novel multi-agent discussion strategy. This strategy integrates self-reflection[13] and Chain-of-Thought (CoT)[17] concepts, as self-reflection focuses on understanding the depth of issues but may lead to rigid thinking. Meanwhile, strategies based on the CoT emphasize the refinement of problems but may lack in-depth analysis of specific issues. Therefore, we combine the two approaches by guiding agents through prompts to decompose evaluation questions and focus on only one sub-question in each discussion round. During each round of the discussion, agents engage in self-reflection, considering peer inputs to enrich issue comprehension. This approach strengthens agents’ self-assessment and critical thinking, broadening their evaluation scope for open-ended text and aligning results more closely with human evaluations.
@@
Furthermore, our framework introduces a feedback mechanism at the end of each discussion round to evaluate the quality and efficiency of the discussions, encouraging agents to reach a consensus. The comprehensive evaluation report generated by our framework details error types, specific locations, in-depth explanations, and corresponding scores.
For practical applications in industry, we provide two report formats: a question-and-answer format for strategy analysis and a text report designed to help business personnel quickly identify errors and facilitate iterative improvement of LLMs. Our framework has achieved significant results in the story text evaluation task at Alipay, markedly enhancing the efficiency of model iteration.
@@
    We propose a novel method to integrate self-reflection and CoT in our multi-agent framework. Additionally, we creatively introduce a feedback mechanism at the end of each discussion round to resolve disagreements and facilitate the achievement of consensus.
@@
    We conduct comprehensive experiments on two English and two Chinese story text datasets, including one constructed based on Alipay’s business story text dataset. Our experimental results showcase the effectiveness of our framework and its high correlation with human evaluations.  [^1] We have made the datasets and results used in our experiments publicly available at https://github.com/kse-ElEvEn/MATEval. Due to the user privacy of Alipay, we cannot make the “Ant” dataset public.
@@ 2 Related Work
LLM-based Evaluators: GPTScore[4] utilizes models such as GPT-3 to evaluate text quality, predicated on the assumption that generative pre-trained models assign higher probabilities to high-quality generated texts by given instructions and context. Recent studies also explore the potential of using ChatGPT as an NLG evaluator[15]. G-Eval[12] demonstrates the evaluation of NLG outputs using prompts in LLMs like ChatGPT through Chain-of-Thought (CoT) methods.
@@ 3.1 Design of the Framework
The framework accepts text as input, which may contain various quality issues. The output is a detailed evaluation report outlining error type, location, explanation, and score. We present the results in two formats: one is a Q&A format conducive to evaluating correlation, allowing easy extraction of correlation scores for similarity calculations. The other is a report format that is conducive to iteration by relevant business personnel. This enables business personnel to quickly identify text issues and refine models using the analysis reports, enhancing efficiency. This is shown in the right part of Figure 1.
@@ 3.2 Application of Different Roles
Figure 2: The diagram includes prompts and dialogue that incorporates a process of discussion with self-reflection, CoT, feedback mechanisms and final summary.
Evaluator Agent: The core element in the framework is the evaluator, for which we use GPT-4 to conduct multi-round evaluations and responses that are guided through carefully designed prompts. The evaluator stores and processes statements from other agents, using this as a reference for dialogue history. Each agent not only receives responses from others but also provides their own statements, with the entire process requiring minimal human intervention.
@@
Summarizer Agent: After all discussions are concluded, the summarizer compiles the entire process and outcomes. It provides a Q&A format evaluation report, detailing the identification, analysis, and scoring of various issues. Additionally, we provide a comprehensive text-based format evaluation report that includes detailed problem descriptions and is easy to read to help improve model performance in industrial production.
@@ 3.4 Combined Self-reflection and CoT
Combined Self-reflection and CoT: As shown in Algorithm 1. Combining self-reflection and the CoT is an important strategy employed in our framework. Agents autonomously decompose the question according to the prompt, focusing on one sub-question in each round of multiple discussions: $\mathcal{D}(\mathcal{Q})=\{\mathcal{Q}_{\text{1}},\mathcal{Q}_{\text{2}},\ldots,\mathcal{Q}_{\text{n}}\}$, $\mathcal{P}_{\text{i}}=\mathcal{I}(\mathcal{Q}_{\text{i}},\mathcal{H})$ , where $\mathcal{I}$ represents the formation of preliminary ideas based on the prompt, $\mathcal{Q}$ is the evaluation task, $\mathcal{Q}_{\text{n}}$ is the sub-question and $\mathcal{H}$ is the history information. They then optimize their statements through self-reflection: $\mathcal{R}_{\text{i}}=\mathcal{S}(\mathcal{P}_{\text{i}},\mathcal{H})$. Next, update the history: =$\mathcal{H}=\mathcal{H}\cup\{\mathcal{R}_{i}\}$. After each round of discussion, a feedback provider evaluates the discussion to reduce repetition and disagreement: $\mathcal{E}=\text{E}(\mathcal{H})$ Finally, a summarizer compiles all statements to produce the evaluation report: $\mathcal{R}_{\text{final}}=\text{Summary}(\mathcal{H})$. The overall prompt and the flow of the discussion are illustrated in Figure 2.
@@ 4.2 Dataset
We mainly apply our framework to the evaluation of story texts generated by LLMs in Alipay business scenarios. So in the experiment, we mainly focus on two open-ended story datasets: ROCStories (ROC)[6] and WritingPrompts (WP)[6]. Considering GPT-4’s context length limitations and the need for storing multi-round discussion contexts, we truncate WP stories to the first 200 words, ensuring textual integrity at the truncation point. To test model generalizability, we conduct similar experiments on two Chinese datasets. These include the Chinese LOng Text understanding and generation (LOT)[7] dataset, comprising human-written stories averaging 106 words, and a dataset of Chinese fairy tales constructed using prompts from Alipay’s business data with GPT-3.5 named Ant, mainly involving fables and fairy tales with an average story length of 125 words. Considering GPT-4’s request rate limits and high usage costs, we select the first 200 stories from each dataset for multi-agent discussion experiments.
@@ 4.3 Compared Methods
Referenced Metrics: The BLEU[14] score is used to evaluate lexical similarity between candidate and reference texts. ROUGE-L[10] focuses on the longest common subsequence to assess the fluency and coherence of texts. And RUBER-BERT[5], an enhancement of the original RUBER model with BERT’s contextual embeddings, includes both referenced and unreferenced versions. The referenced version, RUBER-BERTr, measures the similarity between candidate responses and reference texts using BERT word embeddings.
@@
  • Chain-of-Thought (CoT): Agents break down the assessment problem through prompts, solving one sub-problem in each discussion round.
@@ 4.4 Experimental Results
Our experiments on the ROC and WP datasets are presented in Table 1, using MATEval framework strategies to evaluate narrative texts. We calculate Spearman ($\rho$) and Kendall ($\tau$) correlation coefficients to compare models’ evaluations with human judgments.
Table 1: Correlation of evaluation results with human judgment using different models and different strategies of MATEval on the ROC/WP dataset, where SA stands for Single-Agent, SR denotes Self-Reflection, and CoT represents Chain-of-Thought. The symbols $\rho$/$\tau$ respectively indicate the Spearman/Kendall correlation. The highest correlation values are highlighted in bold.
@@
In analyzing the strategies used within the MATEval framework, we found that employing self-reflection or the CoT independently produces unstable results across different error types. In some cases, these methods even underperformed compared to single-agent evaluations. This might be due to inherent flaws when applying these strategies separately. For example, self-reflection can lead to rigid thinking in multiple discussion rounds, where agents often repeat earlier content without adding new insights. On the other hand, using CoT alone often results in superficial and divergent perspectives, offering only a basic analysis of each error type without delving deeper.
The combination of self-reflection and CoT achieved the best overall correlation, particularly excelling over other methods in evaluating Logical Inconsistency, Discontinuity, and Inappropriate Lexical Choice. It significantly improves the evaluation of Discontinuity compared to the single-agent method, demonstrating the framework’s high sensitivity to textual coherence. However, its effectiveness was lower for Repetition and Factual Error. Agents often misidentified emotionally similar sentences as repetitive, despite clear definitions. Furthermore, the framework’s evaluation of Factual Errors was limited by LLMs constraints, specifically the absence of external knowledge affecting common sense error detection, highlighting the need to integrate external knowledge for future multi-agent framework enhancements.
@@ 4.5 Ablation Study
To assess the effectiveness of different modules in MATEval, we conducted ablation experiments on the ROC dataset, involving the removal of the feedback mechanism, omitting Q&A format explanations, and not using a multi-agent approach. Table 3 demonstrates that the complete MATEval framework surpasses its ablated versions, confirming the significance of both feedback mechanisms, explanations and multi-agent methods. Specifically, this establishes the importance of our feedback mechanisms in promoting discussion consensus and enhancing relevance. It also proves that providing scoring explanations by LLMs significantly enhances evaluation effectiveness.
@@ 4.6 Generalization Experiments
To verify the generalizability of our framework across different languages and in the industrial field, we experimented with two Chinese datasets: LOT and a story text dataset Ant, derived from Alipay’s business data. The findings in Table 2 were similar to those from English datasets. Interestingly, single-agent evaluations often perform better in Chinese, possibly due to its unique language and sentence structure. This suggests that optimizing agents for different languages may require tailored adjustments for optimal performance.
Figure 3: The schematic diagram illustrating the comparison of results generated by the MATEval framework and other methods.
@@ 4.7 Case Study
To showcase the MATEval framework’s effectiveness in real-world industrial settings, we illustrated its application using a story text example shown in Table 3. We compared our framework against manual evaluation, single-agent evaluation, and the strategy-less One_by_One evaluation method. Experimental results demonstrate that our method is basically consistent with human evaluations, unlike other methods, which show some gaps, thereby confirming the high correlation between our approach and human evaluation.
@@ 5 Conclusion
In this paper, we proposed the MATEval framework, which enhances the evaluation performance of open-ended story text generated by LLMs in the industrial field. Extensive experiments show that MATEval’s evaluation results on two classic story datasets are more aligned with human preferences than those of existing methods. In the Alipay industrial scenario, our framework significantly improves review efficiency and evaluation accuracy, serving as an effective aid. [^2] This work was supported by Ant Group [^3] This work was supported by the Natural Science Foundation of China (Grant No. U21A20488). We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper.